51. The dimensions of a metallic cuboid are 100 cm x 80 cm * 64 cm. It is melted and recast into a cube. Find the surface area of the cube.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: Explaination: Dimensions of the metallic cuboid are 100 cm × 80 cm × 64 cm 100 × 80 × 64 = a3 (where a = side of cube) ⇒ 10 × 2 × 4 = a ⇒ a = 80cm Surface area of the cube = 6a² = 6(80)² = 6 × 80 × 80 = 38400 cm²